OpenStudy (anonymous):

Find the four arithmetic means between -21 and -36. I got -24,-27,-30 and -33. Could you please show the work. I am trying to find out where I went wrong in my formula

hartnn (hartnn):

arithmatic mean(AM) between 'a' and 'b' is (a+b)/2 so what is AM between -21 and -36 ??

hartnn (hartnn):

can you find out ?

OpenStudy (anonymous):

Im not sure exactly. We just started working on this today.

hartnn (hartnn):

its not at all difficult. i'll give you an example. AM between -12 and -18 will be (-12-18)/2 = -20/2 = -15. so, in same way , what is AM between -21 and -36 ??? just give it a try.

OpenStudy (anonymous):

a = -21 a6 = -36 a + 5 d = -36 -21 + 5d = -36 5d = -15 d = -3

OpenStudy (sirm3d):

if there are four arithmetic means, there's a term before and a term after. That should be six terms in arithmetic progression.

OpenStudy (anonymous):

Lol..@sirm3d

OpenStudy (sirm3d):

you beat me again. hahaha.